OpenStudy (anonymous):

find the 10th term of the sequence whose 2nd term is 28 and 17th term is -2

OpenStudy (anonymous):

difference = -2 - 28 = -30 over 17 - 2 = 15 terms common difference = -30/15 = -2 10th term = 28 + 8(-2) = 12